# Secure IVSHMEM: End-to-End Shared-Memory Protocol with Hypervisor-CA Handshake and In-Kernel Access Control

### v1

@@ V-C Application Library Implementation
By combining zero-copy ring buffers with doorbell interrupts followed by brief polling, this API achieves near-native IVSHMEM performance while enforcing end-to-end authentication. We evaluate its throughput in Section VI.
VI Measurements
The experiments were run on two Linux guest VMs, each using the 6.12.10-0-lts kernel and provisioned with 2 GiB of RAM and 4 vCPUs. The host is an x86_64 machine with an Intel® Core™ Ultra 7 155H (VT-x enabled, 400 MHz–4.8 GHz) and 32 GiB of DDR memory. To minimize interference, each VM was pinned to its own physical cores, and both the control channel and IVSHMEM devices were instantiated via QEMU’s full-virtualized interfaces
VI-A Latency Overhead for Handshake
@@
TABLE II: Round-trip latency for a single 32-bit integer. Secure IVSHMEM incurs a 150 µs handshake cost, but per-message latency afterward is within 5 % of the vanilla baseline.
@@
As shown in Table II, although the initial handshake adds a modest 150 µs one-time overhead, the steady-state data-plane latency (8.4 µs) is almost identical to vanilla IVSHMEM (8.1 µs). This shows that, although the initial handshake introduces some latency, our security mechanism adds negligible per-message overhead once the session is established.
VI-B Bandwidth Overhead
@@
For each experiment, a total of 32 GiB of random data was transferred to evaluate raw throughput and quantify the overhead introduced by our secure IVSHMEM protocol.
@@
Figure 5: Bandwidth comparison of Secure IVSHMEM Protocol versus Vanilla IVSHMEM.
@@
Figure 5 compares vanilla IVSHMEM with Secure IVSHMEM across message sizes from $2^{6}$ to $2^{15}$ bytes. At small message sizes ($\leq 2^{8}$ B), the added kernel-module hooks introduce up to a 20–25% throughput reduction. However, this overhead quickly vanishes as messages grow. For transfers $\geq 1$ KiB, Secure IVSHMEM’s bandwidth falls within 5 % of the unmodified baseline, showing that the extra hooking logic imposes only a negligible cost at larger message sizes.
VI-C Additional Performance Mitigation
Figure 6: Throughput versus number of producer/consumer pairs for Vanilla IVSHMEM and Secure IVSHMEM with channel-separated mode.
@@ VI-C Additional Performance Mitigation
One of the interesting things is that channel separation brings additional performance gains in multi-producer, multi-consumer scenarios. Figure 6 depicts throughput as a function of the number of concurrent producer/consumer pairs (separate IVSHMEM channels). With a single channel, both vanilla and Secure IVSHMEM sustain $\sim 3.58\,\mathrm{GiB}/\mathrm{s}$. As the channel count increases to eight, per-channel throughput decreases to $\sim 1.56\,\mathrm{GiB}/\mathrm{s}$ for vanilla and $\sim 1.23\,\mathrm{GiB}/\mathrm{s}$ for the secure variant. This performance gain stems from reduced lock contention and parallel, lock-free processing across channels.
Figure 7: Performance benefit of the ring buffer versus alternative read/write on Vanilla IVSHMEM.
Finally, Figure 7 compares the standard read/write interface against our optimized ring-buffer API on vanilla IVSHMEM. At small messages (64 B), the ring buffer yields a modest gain ($\sim 8\%$), but at larger sizes ($\geq 1\,\mathrm{KiB}$) it delivers up to an $80\%$ throughput improvement (e.g., 3.8 GiB/s vs. 2.1 GiB/s at 1 MiB messages). This validates that batching and zero-copy techniques in the ring buffer significantly reduce syscall and copy overhead.
@@
In summary, although our Secure IVSHMEM protocol incurs a one-time handshake latency when establishing each channel, it introduces negligible per-message overhead in both latency and bandwidth once the channel is established; moreover, channel separation combined with a ring-buffer API delivers additional scalability and throughput gains in multi-producer/multi-consumer scenarios by reducing lock contention.
VI-D Security Validation Experiments
@@
Out-of-Bounds Data Access
@@
Control-Section Access Violation
@@
TABLE III: Invalid Access Test Results: all invalid attempts were correctly blocked.
VII Discussion
@@
While we focus here on authentication and integrity, confidentiality could be added via symmetric encryption. A TLS-style key-exchange (for example, ephemeral Diffie–Hellman over the control channel) would introduce only a modest one-time handshake delay and encryption overhead sacrificing zero-copy data-plane performance.
@@
Our protocol and kernel module require applications to link against the IVSHMEM-specific library and invoke its API, rather than using standard socket or networking calls. Achieving full transparency[19][18][17][19]—so that unmodified applications could communicate over IVSHMEM as if it were TCP/IP or any other network transport—is beyond the scope of this work and is left for future exploration (e.g., via socket-API interposition or hypervisor-level redirection).
VIII Conclusion
We have presented Secure IVSHMEM, a protocol that delivers end-to-end authentication and integrity for in-host shared-memory channels without sacrificing zero-copy performance. By treating the hypervisor as a trusted CA and implementing a hypervisor-mediated mutual-authentication handshake, our design prevents spoofing and impersonation attacks, while dynamic kernel-module hooks enforce fine-grained channel access control. Microbenchmarks show that the one-time handshake incurs less than 200 µs latency and that subsequent data transfers achieve near-native IVSHMEM throughput and $\leqslant{5\%}$ round-trip latency overhead. Secure IVSHMEM is therefore well suited for safety-critical, high-performance environments—such as automotive systems—where untrusted services share memory in the same host.

### v2

@@ 5.3. Application Library Implementation
By combining zero-copy ring buffers with doorbell interrupts followed by brief polling, this API achieves near-native IVSHMEM performance while enforcing end-to-end authentication. We evaluate its throughput in Section 6.
6. Measurements
The experiments were run on two Linux guest VMs, each using the 6.12.10-0-lts kernel and provisioned with 2 GiB of RAM and 4 vCPUs. The host is an x86_64 machine with an Intel® Core™ Ultra 7 155H (VT-x enabled, 400 MHz-4.8 GHz) and 32 GiB of DDR memory. To minimize interference, each VM was pinned to its own physical cores, and both the control channel and IVSHMEM devices were instantiated via QEMU’s full-virtualized interfaces
6.1. Latency Overhead for Initial Handshake
@@
Table 2. Round-trip latency for a single 32-bit integer. Secure IVSHMEM incurs a 90µs handshake cost, but per-message latency afterward is within 5 % of the vanilla baseline.
@@
As shown in Table 2, although the initial handshake adds a modest 90µs one-time overhead, the steady-state data-plane latency (8.4 µs) is almost identical to vanilla IVSHMEM (8.1 µs). This shows that, although the initial handshake introduces some latency, our security mechanism adds negligible per-message overhead once the session is established.
6.2. Kernel Module Enforcement Overhead
@@
For each experiment, a total of 32 GiB of random data was transferred to evaluate raw throughput and quantify the overhead introduced by granular access control via a kernel module.
@@
Figure 7. Bandwidth comparison of Secure IVSHMEM Protocol versus Vanilla IVSHMEM.
@@
Figure 7 compares the performance of IVSHMEM with and without access control, across message sizes ranging from $2^{6}$ to $2^{15}$ bytes. For small messages (¡= $2^{8}$ B), the integration of kernel-module hooks incurs a throughput reduction of approximately 20–25%. However, this overhead diminishes rapidly with increasing message size. For transfers ¿= 1 KiB, the bandwidth of Secure IVSHMEM remains within 5% of the unmodified baseline, indicating that the additional hooking logic imposes minimal overhead at larger scales.
In terms of latency, the kernel-module integration introduces an overhead of up to 25–30% for small messages, but this impact also decreases as the message size grows. For messages $\geq$ 1 KiB, the latency overhead drops to below 10%, demonstrating that the cost of access control becomes negligible for larger payloads.
6.3. Performance Optimization
Figure 8. Performance improvements achieved through the use of a ring buffer and anticipation timing window. Channel separation further enhances bandwidth in multi-producer, multi-consumer scenarios
Figure 8 compares the performance of three IVSHMEM communication designs across increasing shared memory sizes: a basic message-passing scheme, a ring-buffer implementation, and a design enhanced by an anticipation timing window.
The results demonstrate substantial performance gains from both optimizations. Applying the ring-buffer architecture—which leverages batching and zero-copy techniques—reduces system call and memory copy overhead and increases bandwidth by more than 30% once the shared-memory size reaches $\geq 2^{10}$ bytes, with an improvement of approximately 46% at $2^{10}$ and sustained gains above 30% at larger sizes. The anticipation timing window further improves throughput by minimizing the cost of frequent interrupt handling, delivering performance boosts of at least 4% across all tested sizes. Together, these enhancements enable more efficient use of shared memory and higher communication bandwidth as the message size increases.
@@ 6.3. Performance Optimization
Additionally, one of the interesting things is that channel separation brings additional performance gains in multi-producer, multi-consumer scenarios. Figure 8 depicts throughput as a function of the number of concurrent producer/consumer pairs (separate IVSHMEM channels). With a single channel, both vanila and Secure IVSHMEM sustain $\sim 3.58\,\mathrm{GiB}/\mathrm{s}$. As the channel count increases to eight, per-channel throughput decreases to $\sim 1.56\,\mathrm{GiB}/\mathrm{s}$ for vanilla and $\sim 1.23\,\mathrm{GiB}/\mathrm{s}$ for the secure variant. This performance gain stems from reduced lock contention and parallel, lock-free processing across channels.
@@
In summary, although our Secure IVSHMEM protocol incurs a one-time handshake latency when establishing each channel, it introduces negligible per-message overhead in both latency and bandwidth once the channel is established; moreover, channel separation combined with a ring-buffer and anticipation timing window delivers additional scalability and throughput gains in multi-producer/multi-consumer scenarios by reducing lock contention.
6.4. Security Validation Experiments
@@
Out-of-Bounds Data Access:
@@
Control-Section Access Violation:
@@
Table 3. Invalid Access Test Results: all invalid attempts were correctly blocked.
7. Discussion
@@
While we focus here on authentication and integrity, confidentiality could be added via symmetric encryption. A TLS-style key-exchange (for example, ephemeral Diffie-Hellman over the control channel) would introduce only a modest one-time handshake delay and encryption overhead sacrificing zero-copy data-plane performance.
@@
Our protocol and kernel module require applications to link against the IVSHMEM-specific library and invoke its API, rather than using standard socket or networking calls. Achieving full transparency(Wang et al., 2008; Zhang and Liu, 2016; Zhang et al., 2007; Wang et al., 2008)—so that unmodified applications could communicate over IVSHMEM as if it were TCP/IP or any other network transport—is beyond the scope of this work and is left for future exploration (e.g., via socket-API interposition or hypervisor-level redirection).
8. Conclusion
We have presented Secure IVSHMEM, a protocol that delivers end-to-end authentication and integrity for in-host shared-memory channels without sacrificing zero-copy performance. By treating the hypervisor as a trusted CA and implementing a hypervisor-mediated mutual-authentication handshake, our design prevents spoofing and impersonation attacks, while dynamic kernel-module hooks enforce fine-grained channel access control. Microbenchmarks show that the one-time handshake incurs less than 100 µs latency and that subsequent data transfers achieve near-native IVSHMEM throughput and $\leq{5\%}$ round-trip latency overhead. Thus, Secure IVSHMEM is well-suited for safety critical, high performance environments such as automotive systems, where untrusted services share memory in the same host.